OpenStudy (anonymous):

find the exact value of the expression given that secx = (3/2), cscy = 3, and x and y are in Quadrant I cos(x - y)

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ cos(\theta)=\cfrac{adjacent}{hypotenuse}\implies \cfrac{a}{c}\qquad sin(\theta)=\cfrac{opposite}{hypotenuse}\implies \cfrac{b}{c}}}\\ \quad \\ \quad \\ sec(x)=\cfrac{3}{2}\qquad csc(y)=3=\cfrac{3}{1}\\ \quad \\ \quad \\ sec(x)=\cfrac{3}{2}\implies \cfrac{1}{cos(x)}=\cfrac{3}{2}\implies {\color{blue}{ \cfrac{2}{3}=cos(x)}}\implies \cfrac{a=2}{c=3}\\ \quad \\ csc(y)=\cfrac{3}{1}\implies \cfrac{1}{sin(y)}=\cfrac{3}{1}\implies {\color{blue}{ \cfrac{1}{3}=sin(y)}}\implies \cfrac{b=1}{c=3}\)

OpenStudy (jdoe0001):

now, for the angle "x", you need the "b" component, that is the opposite side for that you can use the pythagorean theorem \(\bf c^2=a^2+b^2\implies \pm\sqrt{c^2-a^2}={\color{blue}{ b}}\) for the angle "y" you need the "a" component, that is the adjacent side for that you use the pythagorean theorem as well \(\bf c^2=a^2+b^2\implies \pm\sqrt{c^2-b^2}={\color{blue}{ a}}\) notice that the pythagorean theorem uses a radical, which can give a positive or negative value, both valid so which one do we pick? well, we know that "x" and "y" are in Quadrant I, that makes a,b and c, all positive and to expand cos(x-y), you'd need both cosines and sines